# Unbiased Learning to Rank Meets Reality: Lessons from Baidu’s Large-Scale Search Dataset

### v1

@@ Abstract.
Unbiased learning-to-rank (ULTR) is a well-established framework for learning from user clicks, which are often biased by the ranker collecting the data. While theoretically justified and extensively tested in simulation, ULTR techniques lack empirical validation, especially on modern search engines. The dataset released for the WSDM Cup 2023, collected from Baidu’s search engine, offers a rare opportunity to assess the real-world performance of prominent ULTR techniques. Despite multiple submissions during the WSDM Cup 2023 and the subsequent NTCIR ULTRE-2 task, it remains unclear whether the observed improvements stem from applying ULTR or other learning techniques. We revisit and extend the available experiments. We find that unbiased learning-to-rank techniques do not bring clear performance improvements, especially compared to the stark differences brought by the choice of ranking loss and query-document features. Our experiments reveal that ULTR robustly improves click prediction. However, these gains in click prediction do not translate to enhanced ranking performance on expert relevance annotations, implying that conclusions strongly depend on how success is measured in this benchmark.
submissionid: 4266journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington D.C., USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington D.C., USA
@@ 1. Introduction
Historically, the field of learning-to-rank employed human experts to annotate the relevance of search results to create training data for ranking models (Qin and Liu, 2013; Dato et al., 2016, 2022). As expert labeling is expensive (Chapelle and Chang, 2011), infeasible in certain applications (Wang et al., 2016), and potentially misaligned with user preferences (Sanderson et al., 2010), many practitioners seek to leverage implicit user feedback, often collected in the form of clicks. However, clicks are usually a biased signal of relevance, as they often depend on the position of an item (Joachims et al., 2005), its surrounding items (Craswell et al., 2008; Zhuang et al., 2021), or even the user’s trust in the system to place relevant items on top (Joachims et al., 2005; Agarwal et al., 2019a; Vardasbi et al., 2020b). Over the years, the field of unbiased learning-to-rank (ULTR) has proposed methods to mitigate biases when training ranking models on click data.
@@
Figure 1. Original comparison on Baidu-ULTR, showing four ULTR methods not significantly outperforming a naive method not correcting for position bias. Data from (Zou et al., 2022, Table 6) and visualization by us.
@@
Zou et al. (2022) train a MonoBERT ranker and provide baseline results of this model fine-tuned on four standard ULTR methods: inverse propensity scoring (IPS(Wang et al., 2016; Joachims et al., 2017), regression expectation-maximization (RegressionEM(Wang et al., 2018), the dual learning algorithm (DLA(Ai et al., 2018a), and pairwise debiasing (PairD(Hu et al., 2019). Interestingly, Zou et al. (2022) find that these prominent ULTR methods do not perform significantly better than a naive BERT model trained on clicks without any bias correction, [^1] Zou et al. (2022) do find ULTR to be beneficial on frequent head queries compared to infrequent tail queries. as we report in Figure 1. In subsequent work, the Baidu-ULTR dataset was used at multiple ranking competitions (the WSDM Cup 2023 (Zou et al., 2023b, a) and NTCIR’s ULTRE-2 (Niu et al., 2023)), in which multiple teams reported noticeably higher ranking performance, with and without ULTR techniques. It therefore remains unclear whether the observed improvements were due to unbiased learning-to-rank, and ultimately whether ULTR techniques improve performance on this dataset.
@@
    Three competitions have been conducted on the Baidu-ULTR dataset, with participants reporting vastly improved ranking results (although not necessarily using ULTR). A preliminary experiment by us revealed that a random ranking of the annotated test dataset achieves a DCG@10 $\approx 6.69$, substantially higher than all baselines in Figure 1 and thus, hinting at potential problems in the original experiments.
@@
    And lastly, a closer look at the Baidu-ULTR dataset and the original codebase revealed some disputable design decisions. IPS, for example, requires an estimate of position bias on the given dataset, while the creators of the Baidu-ULTR dataset confirmed using a hardcoded position bias from the ULTRA library in their experiments (Tran et al., 2021). We also found that placeholder documents make up for almost $20\%$ of the Baidu-ULTR dataset (see Section 3), with an unknown impact on the trained models.
@@
We answer (RQ1) negatively as we cannot find robust and significant improvements in a fair comparison between ULTR-based and naive methods. We address (RQ2) by reporting minor but existing differences between ULTR methods and by showcasing large and reliable differences due to ranking losses and input features. Regarding (RQ3), we find that while certain ULTR methods improve over their naive counterpart if applied during language model training, the interactions of ULTR and transformer-based models are not well understood and can even lead to decreased performance.
@@ 4. Unbiased Learning-to-Rank Methods
(3) $\begin{split}\mathcal{L}^{\textup{TT}}=\mathcal{L}\left(\sigma\left(\tilde{e}(k)+\tilde{r}(q,d)\right);c\right).\end{split}$
@@
(7) $\mathcal{L}^{\textup{PairD}}=\frac{\mathcal{L}(\tilde{r}(q,d);c)}{\tilde{e}^{+}(k)\tilde{e}^{-}(k)}+\left\|\tilde{e}^{+}(k)\right\|+\left\|\tilde{e}^{-}(k)\right\|.$
@@ 5.1. Training and evaluation procedure
All models are trained on our reranking dataset, and we use a 50/50 random split of our test click dataset (see Table 1) for validation and testing. We evaluate ranking performance on $6,985$ annotated queries, where experts rated each query-document pair’s relevance on a scale from 00 to $4$. We point to Zou et al. (2022) for more details on the annotation process. To remain focused, we do not use available side information such as dwell-time, off-screen scrolling, and returns to the result page. Instead, we use query and document content, document position and click label as our training data.
@@ 5.2.1. Language models
We train cross-encoders on the Baidu-ULTR dataset from scratch in Jax (Bradbury et al., 2018), building on the FlaxBERT implementation from Huggingface. [^10] https://huggingface.co/docs/transformers/model_doc/bert#transformers.FlaxBertForPreTraining The input to our BERT models is: [CLS] query [SEP] doc_title [SEP] doc_abstract. Following previous work on Baidu-ULTR (Zou et al., 2022; Chen et al., 2023b), we truncate the input to a maximum length of 128 tokens. Compared to the original BERT training scheme (Devlin et al., 2019), we keep the masked-language modeling task (masking $30\%$ of input tokens at random) but discard the next-sentence prediction task. Instead, we follow the MonoBERT (Nogueira et al., 2019) setup and apply a linear layer to the BERT CLS token to output a click prediction score. We train multiple models with different click-based loss functions: naive click prediction with a pointwise (binary cross-entropy) and a listwise (softmax cross-entropy) loss, pointwise and listwise IPS, a pointwise two-tower loss, and a listwise DLA loss (see Section 4). All loss functions were built on top of the Rax library (Jagerman et al., 2022).
@@ 5.3. Hyperparameter tuning
We perform extensive hyperparameter tuning for a fair comparison of the reranking models in our experiments. Given the immense combinatorial space of hyperparameters, methods, and datasets, we adopt an incremental hyperparameter tuning strategy as advocated by Godbole et al. (2023). First, we tune our model architecture per dataset based on the pointwise naive model. We tune the number of hidden dimensions $\in\{64,128,256,512,1024\}$ and the number of layers $\in\{2,3,4,5\}$ over three random seeds. As both network depth and width can interact with the learning rate, we tune each parameter combination over three learning rates $\in\{0.001,0.0005,0.0001\}$, i.e., we treat the learning rate as a nuisance parameter (Godbole et al., 2023). We adopt this incremental tuning procedure as we found no major discrepancies in model architecture between ULTR methods but instead between datasets. Subsequently, we adopt the architecture of the pointwise naive model per dataset for all other methods.
A five-layer perceptron with 512 hidden dimensions yields optimal results for our LTR features, while a five-layer perceptron with 256 dimensions performs best for the Baidu BERT embeddings. Additionally, a two-layer perceptron with 256 dimensions was the most suitable choice for our BERT embeddings. Given each base model architecture, we tune the final dropout $\in\{0,0.3\}$ and learning rate $\in\{0.001,0.0005,0.0003,0.0001\}$ for each method and dataset over three random seeds. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\varepsilon=1e^{-8}$. Most models work best with a learning rate of $0.0001$ and do not benefit from dropout. The full list of hyperparameters is available in our online repository. Given the final hyperparameters for each method, we train all methods for 50 epochs, stopping early after five epochs of no improvement of the validation loss computed on clicks. Lastly, we highlight that while this tuning step improved the performance of all methods, our main findings are consistent across many different hyperparameter combinations.
@@ 6. Results
Figure 3. Comparing ULTR method on pre-trained BERT embeddings and LTR features. We display the average ranking performance measured in DCG@10 over five independent runs and plot a bootstrapped 95% confidence interval. The grey dotted line indicates the performance of a random ranker.
@@ 6.1. Unbiased learning-to-rank yields no or tiny improvements on Baidu-ULTR
Second, given a set of input features, the choice of ranking loss used as a base for ULTR methods matters greatly (comparing the colored groups in Figure 3). Methods based on LambdaRank outperform methods based on the listwise softmax loss, which outperform those based on the pointwise binary cross-entropy loss. In contrast, applying ULTR techniques yield marginal improvements at most compared to their naive, non-debiasing counterpart.
@@
Overall, we can answer research questions (RQ1) and (RQ2): on the Baidu-ULTR dataset, unbiased learning-to-rank techniques do not bring any clear benefits, particularly when contrasted with the significant and reliable differences based on the choice of query-document features and ranking loss. Our reranking results confirm the initial findings of Zou et al. (2022).
@@ 6.2. Language model training is sensitive to the choice of unbiased learning-to-rank method
Table 2 gives an overview of applying ULTR methods during language model training. In contrast to the inefficacy of ULTR-based rerankers, we observe stark differences due to applying ULTR methods when training language models. Similar to the reranking task, approaches based on the listwise softmax loss perform better than those based on pointwise binary cross-entropy. The additional application of ULTR brings considerable improvements with the pointwise two-tower objective. However, we also observe substantial degradations with both IPS and DLA. The best method is the naive listwise softmax loss without any debiasing objective.
@@ 6.3. Click prediction does not imply ranking performance on annotations
In addition to evaluating ranking performance, we display the negative log-likelihood (NLL) of predicted click scores in Figure 4. We restrict this evaluation to pointwise ULTR methods, making well-defined click predictions. All ULTR methods robustly improve click prediction compared to a naive loss, with the two-tower model showing the largest improvement on all three datasets. Moreover, as discussed in Section 3.2, the propensities learned by RegressionEM are close to those discovered through intervention harvesting and show a strong position bias. This finding suggests that unbiased learning-to-rank robustly captures position bias and can better predict user clicks because of it. Yet, this improved click prediction does not translate to enhanced ranking performance on expert annotations.
@@
Figure 4. Click prediction performance of pointwise methods measured in negative log-likelihood, lower is better.
@@ 7. Discussion
Pointwise Two-Tower $7.456$ $0.217$
Pointwise IPS $6.296$ $0.317$
@@ 7.1. Potential reasons for the failure of ULTR
More complex user behavior. Another hypothesis, initially suggested by Zou et al. (2022), is that the actual user model is more complicated than the simple PBM we consider in this work. This possibility might explain why listwise approaches bring such considerable improvements over pointwise methods and would not be surprising as user studies have identified more complex biases (Agarwal et al., 2019a; Zhuang et al., 2021; Sarvi et al., 2023). However, we point out that ULTR methods bringing no benefits would mean that the substantial position bias we identified is negligible against other types of biases. This assumption, in turn, would contradict related work where PBM-based models trained on more complex user behaviors still improved performance, albeit by not as much as the correct user model (Vardasbi et al., 2020a; Deffayet et al., 2023b).
@@
User-annotator disagreement. Finally, the divergence of click prediction and ranking performance hints at a deeper issue: annotators may not find the same documents relevant as users scrolling through a result page, which might be personalized and part of a multi-query session to find actionable information for the current user needs. We hope more user studies or parallel datasets of clicks and expert annotations can clarify this hypothesis.
@@ 7.3. Implications for the field
Our results starkly contrast with many studies in semi-synthetic simulation (Joachims et al., 2017; Ai et al., 2021; Tran et al., 2021; Saito et al., 2020; Ai et al., 2018a; Hu et al., 2019). They call for adjusting these semi-synthetic experimental setups to better reflect real-world challenges. The richness of this dataset allows for seeding simulations with more plausible data and, therefore, might bridge the gap between simulations and reality. Moreover, the prevalence of transformer-based models and text embeddings in current IR research encourages exploring the interaction of ULTR methods with such models. As shown in Section 6.2, the same techniques applied during language model training, instead of on a small re-ranking model, can yield vastly different results.

### v2

@@ Abstract.
Unbiased learning-to-rank (ULTR) is a well-established framework for learning from user clicks, which are often biased by the ranker collecting the data. While theoretically justified and extensively tested in simulation, ULTR techniques lack empirical validation, especially on modern search engines. The Baidu-ULTR dataset released for the WSDM Cup 2023, collected from Baidu’s search engine, offers a rare opportunity to assess the real-world performance of prominent ULTR techniques. Despite multiple submissions during the WSDM Cup 2023 and the subsequent NTCIR ULTRE-2 task, it remains unclear whether the observed improvements stem from applying ULTR or other learning techniques.
In this work, we revisit and extend the available experiments on the Baidu-ULTR dataset. We find that standard unbiased learning-to-rank techniques robustly improve click predictions but struggle to consistently improve ranking performance, especially considering the stark differences obtained by choice of ranking loss and query-document features. Our experiments reveal that gains in click prediction do not necessarily translate to enhanced ranking performance on expert relevance annotations, implying that conclusions strongly depend on how success is measured in this benchmark.
Learning to rank, Counterfactual learning-to-rank, Click models
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657892isbn: 979-8-4007-0431-4/24/07
@@ 1. Introduction
Historically, the field of learning-to-rank employed human experts to annotate the relevance of search results to create training data for ranking models (Qin and Liu, 2013; Dato et al., 2016, 2022). As expert labeling is expensive (Chapelle and Chang, 2011), infeasible in certain applications (Wang et al., 2016), and potentially misaligned with user preferences (Sanderson et al., 2010), many practitioners seek to leverage implicit user feedback, often collected in the form of clicks. However, clicks are usually a biased signal of relevance, as they often depend on the position of an item (Joachims et al., 2005), its surrounding items (Craswell et al., 2008; Zhuang et al., 2021), or even the user’s trust in the system to place relevant items on top (Joachims et al., 2005; Agarwal et al., 2019a; Vardasbi et al., 2020b). Over the years, the field of unbiased learning-to-rank (ULTR) has proposed methods to mitigate biases when training ranking models on click data (Gupta et al., 2024).
@@
Figure 1. The original experiments on Baidu-ULTR show that none of the four compared ULTR methods outperform a naive method not correcting for position bias. Data from (Zou et al., 2022, Table 6) and visualization by us.
@@
Zou et al. (2022) train a MonoBERT ranker and provide baseline results of this model fine-tuned on four standard ULTR methods: inverse propensity scoring (IPS(Wang et al., 2016; Joachims et al., 2017), regression expectation-maximization (RegressionEM(Wang et al., 2018), the dual learning algorithm (DLA(Ai et al., 2018a), and pairwise debiasing (PairD(Hu et al., 2019). Interestingly, Zou et al. (2022) find that these prominent ULTR methods overall do not perform significantly better than a naive BERT model trained on clicks without any bias correction, [^1] Zou et al. (2022) find that DLA outperforms a naive baseline on frequent head queries. as we report in Figure 1. In subsequent work, the Baidu-ULTR dataset was used at multiple ranking competitions (the WSDM Cup 2023 (Zou et al., 2023b, a) and NTCIR’s ULTRE-2 (Niu et al., 2023)), in which multiple teams reported noticeably higher ranking performance, with and without ULTR techniques. It therefore remains unclear whether the observed improvements were due to unbiased learning-to-rank, and ultimately whether ULTR techniques improve performance on this dataset.
@@
    Three competitions have been conducted on the Baidu-ULTR dataset, with participants reporting vastly improved ranking results, though not necessarily using ULTR. A preliminary experiment by us revealed that a random ranking of the annotated test dataset achieves a DCG@10 $\approx 6.69$, which is higher than all baselines in Figure 1 and suggests potential issues with the original experiments.
@@
    And lastly, a closer look at the Baidu-ULTR dataset and the original codebase revealed some disputable design decisions. IPS, for example, requires an estimate of position bias on the given dataset, while the creators of the Baidu-ULTR dataset confirmed using a hardcoded position bias from the ULTRA library in their experiments (Tran et al., 2021). We also found that placeholder documents make up almost $20\%$ of the Baidu-ULTR dataset (see Section 3), with an unknown impact on the trained models.
@@
We answer (RQ1) negatively as we cannot find robust improvements in ranking performance in a fair comparison between ULTR-based and naive methods. For (RQ2), we report minor but existing differences between ULTR methods and showcase large and reliable differences due to ranking losses and input features. Regarding (RQ3), we find that while certain ULTR methods improve over their naive counterpart if applied during language model training, the interactions of ULTR and transformer-based models are not well understood and can even lead to decreased performance.
@@ 4. Unbiased Learning-to-Rank Methods
(3) $\begin{split}\mathcal{L}^{\textup{TwoTower}}=\mathcal{L}\left(\sigma\left(\tilde{e}(k)+\tilde{r}(q,d)\right);c\right).\end{split}$
@@
(7) $\mathcal{L}^{\textup{PairD}}=\frac{\mathcal{L}(\tilde{r}(q,d);c)}{\tilde{e}^{+}(k)\tilde{e}^{-}(k)}+\|\tilde{e}^{+}\|_{1}+\|\tilde{e}^{-}\|_{1}.$
@@ 5.1. Training and evaluation procedure
All models are trained on our reranking dataset, and we use a 50/50 random split of our test click dataset (see Table 1) for validation and testing. We evaluate ranking performance on $6,985$ annotated queries, where experts rated each query-document pair’s relevance on a scale from 00 to $4$. We point to Zou et al. (2022) for more details on the annotation process. In this work, we do not use available side information such as dwell-time, off-screen scrolling, and returns to the result page. Instead, we focus on query and document content, document position and click label as our training data.
@@ 5.2.1. Language models
We train cross-encoders on the Baidu-ULTR dataset from scratch in Jax (Bradbury et al., 2018), building on the FlaxBERT implementation from Huggingface. [^10] https://huggingface.co/docs/transformers/model_doc/bert#transformers.FlaxBertForPreTraining The input to our BERT models is: [CLS] query [SEP] doc_title [SEP] doc_abstract. Following (Zou et al., 2022; Chen et al., 2023b), we truncate the input to a maximum length of 128 tokens. Compared to the original BERT training scheme (Devlin et al., 2019), we keep the masked-language modeling task (masking $30\%$ of input tokens at random) but discard the next-sentence prediction task. Instead, we follow the MonoBERT (Nogueira et al., 2019) setup and apply a linear layer to the BERT CLS token to output a click prediction score. We train multiple models with different click-based loss functions: naive click prediction with a pointwise (binary cross-entropy) and a listwise (softmax cross-entropy) loss, pointwise and listwise IPS, a pointwise two-tower loss, and a listwise DLA loss (see Section 4). All loss functions were built on top of the Rax library (Jagerman et al., 2022).
@@ 5.3. Hyperparameter tuning
We perform extensive hyperparameter tuning for a fair comparison of the reranking models in our experiments. Given the immense combinatorial space of hyperparameters, methods, and datasets, we adopt an incremental hyperparameter tuning strategy as advocated by Godbole et al. (2023). First, we tune our model architecture per set of input features based on the pointwise naive model. We tune the number of hidden dimensions $\in\{64,128,256,512,1024\}$ and the number of layers $\in\{2,3,4,5\}$ over three random seeds. As both network depth and width can interact with the learning rate, we tune each parameter combination over three learning rates $\in\{0.001,0.0005,0.0001\}$, i.e., we treat the learning rate as a nuisance parameter (Godbole et al., 2023). We adopt this incremental tuning procedure as we found no major discrepancies in model architecture between ULTR methods but instead between sets of input features. Subsequently, we adopt the architecture of the pointwise naive model per dataset for all other methods.
A five-layer perceptron with 512 hidden dimensions yields optimal results for our LTR features, while a five-layer perceptron with 256 dimensions performs best for the Baidu BERT embeddings. Additionally, a two-layer perceptron with 256 dimensions was the most suitable choice for our BERT embeddings. Given each base model architecture, we tune the final dropout $\in\{0,0.3\}$ and learning rate $\in\{0.001,0.0005,0.0003,0.0001\}$ for each method and dataset over three random seeds. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\varepsilon=1e^{-8}$. Most models work best with a learning rate of $0.0001$ and do not benefit from dropout. The full list of hyperparameters is available in our online repository. Given the final hyperparameters for each method, we train all methods for 50 epochs, stopping early after five epochs of no improvement of the validation loss computed on clicks. Lastly, while this tuning step improved the performance of all methods, our main findings are consistent across many different hyperparameter combinations.
@@ 6. Results
Figure 3. Comparing ULTR methods on pre-trained BERT embeddings and LTR features. We display the average ranking performance measured in DCG@10 over five independent runs and plot a bootstrapped 95% confidence interval. The grey dotted line indicates the performance of a random ranker.
@@ 6.1. Unbiased learning-to-rank yields no or tiny improvements on Baidu-ULTR
Second, given a set of input features, the choice of ranking loss used as a base for ULTR methods matters greatly (comparing the colored groups in Figure 3). Methods based on LambdaRank outperform methods based on the listwise softmax loss, which outperform those based on the pointwise binary cross-entropy loss. In contrast, applying ULTR techniques yields marginal ranking improvements at most compared to their naive, non-debiasing counterpart.
@@
Overall, we can answer research questions (RQ1) and (RQ2): on the Baidu-ULTR dataset, unbiased learning-to-rank methods do not consistently improve ranking performance on expert annotations, particularly when contrasted with the significant and reliable differences based on the choice of query-document features and ranking loss. Our reranking results confirm the findings of Zou et al. (2022).
@@ 6.2. Language model training is sensitive to the choice of unbiased learning-to-rank method
Table 2 gives an overview of applying ULTR methods during language model training. In contrast to the inefficacy of ULTR-based rerankers, we observe stark differences when applying ULTR methods during language model training. Similar to the reranking task, approaches based on the listwise softmax loss perform better than those based on pointwise binary cross-entropy. The additional application of ULTR brings considerable improvements with the pointwise two-tower objective. However, we also observe substantial degradations with both IPS and DLA. The best method is the naive listwise softmax loss without any debiasing objective.
@@ 6.3. Click prediction does not imply ranking performance on annotations
In addition to evaluating ranking performance, we display the negative log-likelihood (NLL) of predicted click scores in Figure 4. We restrict this evaluation to pointwise ULTR reranking methods, making well-defined click predictions. All ULTR methods robustly improve click prediction compared to a naive loss, with the two-tower model showing the largest improvement on all three datasets. Moreover, as discussed in Section 3.2, the propensities learned by RegressionEM are close to those discovered through intervention harvesting and show a strong position bias. This finding suggests that unbiased learning-to-rank robustly captures position bias and can better predict user clicks because of it. Yet, this improved click prediction does not translate to enhanced ranking performance on expert annotations.
@@
Figure 4. Click prediction performance of pointwise methods measured in negative log-likelihood; lower is better.
@@ 7. Discussion
Pointwise Two-Tower $7.456$ $0.218$
Pointwise IPS $6.296$ $0.222$
@@ 7.1. Potential reasons for the failure of ULTR
More complex user behavior. Zou et al. (2022) suggest that the actual user model might be more complicated than the PBM considered in this work. This hypothesis might explain why listwise approaches bring such considerable improvements over pointwise methods and would not be surprising as user studies have identified more complex biases (Agarwal et al., 2019a; Zhuang et al., 2021; Sarvi et al., 2023). However, ULTR methods bringing no benefits would mean that the substantial position bias we identified is negligible against other types of biases. This assumption, in turn, would contradict related work where PBM-based models trained on more complex user behaviors still improved performance, albeit by not as much as the correct user model (Vardasbi et al., 2020a; Deffayet et al., 2023b).
@@
User-annotator disagreement. Finally, the divergence of click prediction and ranking performance hints at a deeper issue: annotators may not find the same documents relevant as users scrolling through a result page, which might be personalized and part of a multi-query session to find actionable information for the current user needs. We hope further analysis of Baidu-ULTR or parallel datasets of clicks and expert annotations can clarify this hypothesis.
@@ 7.3. Implications for the field
Our results starkly contrast with many studies in semi-synthetic simulation (Joachims et al., 2017; Ai et al., 2021; Tran et al., 2021; Saito et al., 2020; Ai et al., 2018a; Hu et al., 2019) and call for adjusting semi-synthetic experimental setups to reflect real-world challenges better. The richness of this dataset allows for seeding simulations with more plausible data and, therefore, might bridge the gap between simulations and reality. Moreover, the prevalence of transformer-based models and text embeddings in current IR research encourages exploring the interaction of ULTR methods with such models. As shown in Section 6.2, the same techniques applied during language model training, instead of on a small re-ranking model, can yield vastly different results.